Suppose the demand curve for a product is given by Q = 500 − 2P + 4I, where I is average income measured in thousands of dollars. The supply curve is Q = 3P − 80. If I = 25, find the market-clearing price and quantity for the product. 


1
Expert's answer
2021-03-04T10:40:57-0500

Let's first calculate the demand curve for a product if I=25:


"Q_d=500-2P+4\cdot25=600-2P."

The market equilibrium occurs when "Q_d=Q_s", then:


"600-2P=3P-80,""680=5P,""P=\$136."

Therefore, the market-clearing price for the product is $136.

Finally, substituting "P" into "Q_d" we can find quantity for the product:


"Q=600-2\cdot\$136=328."